I am having such a hard time please help me? the length of a rectangle is 5 cm more than 2 times its width. if the area of the rectangle is 68cm^2, find the dimensions of the rectangle to the nearest thousandth
Answer by vertciel(183)

We don't know the numbers for the length and width, so let's use variables.
Let w be the width.
Length = 5 + 2W
Width = w
Area of rectangle = LW
(5 + 2W)W = 68^2
2W^2 + 5W = 4624
2W^2 + 5W - 4624 = 0
Now, we see that this is a quadratic equation. You can try to factor this equation or just use the quadratic formula.
If you need any more help, let me know.